OpenStudy (anonymous):

Given: F = {(0, 1), (2, 4), (4, 6), (6, 8)} and G = {(2, 5), (4, 7), (5, 8), (6, 9), (7, 5)} (F - G) (6) =

OpenStudy (anonymous):

Try using the law that states (f-g)(6)= f(6)-g(6). Does this help?

OpenStudy (anonymous):

-3 -1 3 it's got to be one of these answers

OpenStudy (anonymous):

Well f(6)-g(6)= 8-9=-1. All you have to do is look at the x-values of f that are equal to 6 and then write down the corresponding y values for that x to get it and the same for the g values.